Two boats are sailing directly toward each other at different speeds. One moves at 20m/s while the other at 32m/s. The boats start 2,500 m apart. How much time does it take before the boats collide?
1
Expert's answer
2018-01-29T10:17:08-0500
The velocity of the boats relative to each other
v=v_1+v_2=20+32=52 m/s
So, the time when the boats would collide
t=L/v=2500/52=48 s